OpenStudy (anonymous):

Distribute: 3(4x + 8)

OpenStudy (kaylaprincess):



Nnesha (nnesha):

multiply both terms in the parentheses by outside number \[\huge\rm a(b+c)= a \times b +a \times c =ab+ac\]